# Privacy Risks and Preservation Methods in Explainable Artificial Intelligence: A Scoping Review

### v2

@@ 5.6 Other techniques
Guarded counterfactuals Training data, model Example-based (counterfactuals) An & Cao (2024)
@@
Interpretation-Nets Training data Interpretable (decision trees) Marton et al. (2024)
@@ 6.5 Prevent direct exposure
Explanations generated by XAI tools should not disclose personally identifiable and/or sensitive information to unauthorized individuals (Chen et al., 2022a). In Use Case 2, when the customer seeks an explanation on his/her application outcome, the explanation might indicate the failure to meet respective thresholds of \$40K for salary and 50 years for age. Revealing actual values of protected features would breach the customer’s privacy when accessed by other actors, such as the bank executive during the customer’s consultation. The customer may, however, subsequently provide consent to the executive to retrieve their personal and financial details from the bank’s records for consultation.
@@ 6.7 Access control of explanations
The content of explanations should be accessible only to authorized users (Blanco-Justicia et al., 2020; Kuppa & Le-Khac, 2021) with details provided on need-to-know basis. In Use Case 2, the customer is authorized to access his/her own explanation as it pertains to their specific application outcome. The bank executive should be authorized to access the explanation only if a human intermediary is required to enhance the process of explanation for the customer. In Use Case 3, the AI developer and quality engineer should be permitted to access explanations and outcomes only for synthetic queries generated to simulate specific error cases rather than for real production data.
@@ 6.8 Upholding of explanation quality
The quality of explanations should not be compromised by the introduction of privacy preservation measures. Explanations must remain useful and meaningful (Shokri et al., 2021) to target stakeholders. In Use Cases 2 and 3, the details contained in the explanations to respective users should assist them in completing their tasks effectively and/or help them interpret the outcome of the AI system.
@@ 6.9 Acceptable run time
The run time of XAI methods, being an important evaluation metric (Bodria et al., 2023), should not deteriorate by introduction of privacy preservation measures. In Use Cases 2 and 3, the explanation recipients should see the outputs within an acceptable timeframe. A long turnaround time may lead to the explanations becoming ineffective for the task at hand.
@@ 6.10 Compliance with applicable AI/privacy regulations
XAI being an AI system, should comply with applicable AI and privacy regulations specfic to the jurisdiction in which it operates. For instance, if the XAI is deployed in Canada with Canadian residents as the target users, it must adhere to the provisions of the Artificial Intelligence and Data Act (AIDA, 2022). Users should be clearly informed of the XAI capabilities of the system including the types and content of explanations and the third parties with whom the explanations may be shared. Furthermore, appropriate consent must be obtained from users, as required by applicable laws and regulations.
@@ 7.1 Summary and implications
The comprehensive review of existing literature facilitated a synthesis of current knowledge on the conflict between privacy and explainability, both being important pillars of Trustworthy AI. The analyzed studies demonstrate that the additional information provided in the form of explanations can benefit adversaries in launching privacy attacks in XAI. We identified and categorized certain types of privacy leakage due to malicious intent of adversaries as intentional causes. These include membership inference, model inversion and model extraction, all of which are demonstrated on explanations generated using different methods. These attacks pose a threat to the privacy of individuals whose data is contained in the training set, thus increasing the risk of identification of individuals or exposure of their sensitive information. Moreover, reconstruction of data from inference-time queries renders active XAI users vulnerable to similar privacy breaches (Zhao et al., 2023; 2021). The threat of model extraction through explanations, targets the confidentiality and intellectual property of model owners. While property inference can expose sensitive aggregates or group properties of the training data, our review found no evidence of such attacks targeting XAI systems specifically. Beyond privacy attacks, ML models exhibit inherent privacy vulnerabilities, such as memorization of training data or overfitting, which can lead to various inferencing attacks. These privacy problems are inherited in XAI systems, and we have categorized them as unintentional causes. Additionally, the explanation content can be at direct threat of privacy breaches by unauthorized users due to lack of access control, or through proxies and correlated fields. Such vectors further compound the privacy risks associated with deploying XAI technologies in sensitive domains.
@@ 7.3 Open issues and challenges
  • Lack of user-centricity in development of privacy preserving XAI: End-users are an integral and inseparable component of XAI as they directly engage and draw insights from the content generated by these systems. While current XAI methods are predominantly model-centric, focussing on model development, evaluation and audit processes (Kaplan et al., 2024), there is an increasing need for a shift towards a user-centric approach. When considering privacy as a system requirement or using existing privacy preserving solutions with explainability methods, human users are absent from the design and development of these approaches.
@@
  • Lack of standardised approach for privacy evaluation of XAI: Existing XAI literature provides evaluation metrics to assess aspects such as faithfulness, complexity, localisation, randomisation and robustness (Hedström et al., 2023). However, though privacy is fundamental in Trustworthy AI, currently there is a lack of standardised approach to evaluate the privacy of explanations. The lack of suitable quantitative metrics in this area prevents an evaluaton of safety with respect to privacy or comparison between methods for selection of those that satisfy the safety requirements of an application domain.
@@
  • Trade-offs in privacy, explainability and utility: The introduction of privacy enhancing technologies often result in adverse effects on model utility (Ezzeddine et al., 2024; Harder et al., 2020) and the quality of explanations. For instance, the perturbation of classifier weights of support vector machines for privacy preservation is observed to deteriorate the classification accuracy and credibility of counterfactuals (Mochaourab et al., 2023). Similarly, the use of differential privacy in neural network models is found to lower its accuracy (Blanco-Justicia et al., 2023) and explanation quality (Liu et al., 2024). Thus when targeting these 3 properties simultaneously in an XAI system, there are often trade-offs involved.
@@
  • Lack of XAI methods that are privacy preserving by design: As emphasized by Hoepman (2014), privacy is a core property of computer systems that requires addressing from system design phase, rather than treated as an add-on. However, current explainability approaches do not consider privacy as a system requirement. Due to the overlook of this safety aspect, design flaws may be introduced which can lead to unintentional privacy leakage. Adversaries may exploit these flaws to cause intentional leakage.
@@
  • Lack of privacy preserving XAI for Gen-AI and LLMs: XAI research has mainly focused on discriminative models that produce decision boundaries, and there has been limited work on developing explainability methods for Gen-AI and LLMs (Schneider, 2024b; Sun et al., 2022; Weisz et al., 2023). Due to the complex structure and vast number of parameters in these models, traditional explainability methods become impractical to them (Zhao et al., 2024). These models have privacy issues, such as memorization of training data that escalates as the models become larger (Carlini et al., 2021). In addition, downstream private datasets used for in-context learning in LLMs, are found to be susceptible to membership inference (Duan et al., 2023). When creating explainability approaches for these systems, privacy should be considered as an integral requirement.
@@ 7.4 Recommendations for future work
  • Improving usability of privacy preserving XAI: Explanations should be designed to meet the diverse information needs of users and integrate user-centric design principles (Ali et al., 2023; Kaplan et al., 2024). The delivery of explanations is required to be in a format that is accessible and meaningful, taking into consideration the varying levels of expertise and requirements of different user groups. This transition would prioritize providing need-to-know information to end-users based on their specific roles within the system. Furthermore, to enhance the effectiveness of explanations, appropriate tools such as interactivity and visualization should be used to enhance the process of explanation and deepen users’ understanding (Bo et al., 2024). In addition, application of the 3C-principle of context, content, and consent (Brunotte et al., 2023) can improve the usability of XAI tools by satisfying their requirements and expectations.
@@
  • Balancing trade-off in privacy, explainability and utility: Determining the appropriate trade-off between the triad of privacy, explainability and utility, can help to achieve the right measure of balance between these properties. This could be achieved with the help of tools or tuning mechanisms. For instance, the use of compatibility matrix (Abbasi et al., 2024) or hyperparameters, such as the privacy budget $\epsilon$ in differential privacy or the parameter k in k-anonymity, is useful in tuning the desired level of these required properties. Similar tuning mechanisms can be integrated in other privacy preserving approaches to achieve the required trade-off. Metrics, such as trade-off score (Abbasi et al., 2024), could be useful to quantify and monitor the balance of these properties enabling researchers and practitioners to adjust the parameters based on the trade-offs involved.
@@
  • Development of XAI methods that are privacy preserving by design: The characteristics of privacy preserving XAI outlined in this review, can aid researchers and developers in building explainability methods that are privacy-enhanced by design (Bozorgpanah & Torra, 2024). Furthermore, there is growing interest in neuro-symbolic approaches (Hitzler et al., 2022) and semantic technologies (Seeliger et al., 2019) as potential solutions as explainable-by-design strategies. Researchers can investigate the privacy safety of these techniques and integrate approaches to prioritize privacy.
  • Privacy preserving XAI for Gen-AI and LLMs: XAI plays a vital role in fostering trustworthiness (Wang & Ding, 2024) and ensuring ethical applications of Gen-AI and LLMs (Luo & Specia, 2024). However, as explainability is introduced in these models, it is important to ensure that it does not exacerbate the inherent privacy issues of these models or create new vulnerabilities. A privacy analysis of explainability methods in the early stages of development and the use of privacy attacks for auditing (Carlini et al., 2021), can boost the development of privacy-enhanced systems. Recently, Chain-of-Thought (CoT) (Wei et al., 2022) prompting has gained traction in eliciting the step-by-step reasoning of models and providing self-explanations in the process. However, the interaction with untrusted LLMs can be a threat to users’ privacy, hence future work can be directed towards privacy preserved mechanisms (Bae et al., 2025) in these models. Methods such as retrieval-augmented generation (RAG) for fine tuning of outputs by augmenting external data sources (Zeng et al., 2024) and mechanistic interpretability (Section 2.3.8) can also be investigated as possible solutions towards achieving privacy preservation XAI. Thus, with the growing accessibility and widespread use of Gen-AI and LLMs, developing appropriate user-centric privacy preserving explainability techniques is an important avenue for further research.

### v3

@@ 5.6 Other techniques
Guarded counterfactuals Training data, model Example-based (counterfactuals) anCounterfactualExplanationWill2024
@@
Interpretation-Nets Training data Interpretable (decision trees) martonExplainingNeuralNetworks2024
@@ 6.5 Prevent direct exposure
Explanations generated by XAI tools should not disclose personally identifiable and/or sensitive information to unauthorized individuals (chenAchievingTransparencyReport2022). In Use Case 2, when the customer seeks an explanation on his/her application outcome, the explanation might indicate the failure to meet respective thresholds of \$40K for salary and 50 years for age. Revealing actual values of protected features would breach the customer’s privacy when accessed by other actors, such as the bank executive during the customer’s consultation. The customer may, however, subsequently provide consent to the executive to retrieve their personal and financial details from the bank’s records for consultation.
@@ 6.7 Access control of explanations
The content of explanations should be accessible only to authorized users (blanco-justiciaMachineLearningExplainability2020; kuppaAdversarialXAIMethods2021) with details provided on need-to-know basis. In Use Case 2, the customer is authorized to access his/her own explanation as it pertains to their specific application outcome. The bank executive should be authorized to access the explanation only if a human intermediary is required to enhance the process of explanation for the customer. In Use Case 3, the AI developer and quality engineer should be permitted to access explanations and outcomes only for synthetic queries generated to simulate specific error cases rather than for real production data.
@@ 6.8 Upholding of explanation quality
The quality of explanations should not be compromised by the introduction of privacy preservation measures. Explanations must remain useful and meaningful (shokriPrivacyRisksModel2021) to target stakeholders. In Use Cases 2 and 3, the details contained in the explanations to respective users should assist them in completing their tasks effectively and/or help them interpret the outcome of the AI system.
@@ 6.9 Acceptable run time
The run time of XAI methods, being an important evaluation metric (bodriaBenchmarkingSurveyExplanation2023), should not deteriorate by introduction of privacy preservation measures. In Use Cases 2 and 3, the explanation recipients should see the outputs within an acceptable timeframe. A long turnaround time may lead to the explanations becoming ineffective for the task at hand.
@@ 6.10 Compliance with applicable AI/privacy regulations
XAI being an AI system, should comply with applicable AI and privacy regulations specfic to the jurisdiction in which it operates. For instance, if the XAI is deployed in Canada with Canadian residents as the target users, it must adhere to the provisions of the Artificial Intelligence and Data Act (aidaActEnactConsumer2022). Users should be clearly informed of the XAI capabilities of the system including the types and content of explanations and the third parties with whom the explanations may be shared. Furthermore, appropriate consent must be obtained from users, as required by applicable laws and regulations.
@@ 7.1 Summary and implications
The comprehensive review of existing literature facilitated a synthesis of current knowledge on the conflict between privacy and explainability, both being important pillars of Trustworthy AI. The analyzed studies demonstrate that the additional information provided in the form of explanations can benefit adversaries in launching privacy attacks in XAI. We identified and categorized certain types of privacy leakage due to malicious intent of adversaries as intentional causes. These include membership inference, model inversion and model extraction, all of which are demonstrated on explanations generated using different methods. These attacks pose a threat to the privacy of individuals whose data is contained in the training set, thus increasing the risk of identification of individuals or exposure of their sensitive information. Moreover, reconstruction of data from inference-time queries renders active XAI users vulnerable to similar privacy breaches (zhaoEfficientPrivacypreservingTreebased2023; zhaoExploitingExplanationsModel2021). The threat of model extraction through explanations, targets the confidentiality and intellectual property of model owners. While property inference can expose sensitive aggregates or group properties of the training data, our review found no evidence of such attacks targeting XAI systems specifically. Beyond privacy attacks, ML models exhibit inherent privacy vulnerabilities, such as memorization of training data or overfitting, which can lead to various inferencing attacks. These privacy problems are inherited in XAI systems, and we have categorized them as unintentional causes. Additionally, the explanation content can be at direct threat of privacy breaches by unauthorized users due to lack of access control, or through proxies and correlated fields. Such vectors further compound the privacy risks associated with deploying XAI technologies in sensitive domains.
@@ 7.3 Open issues and challenges
  • Lack of user-centricity in development of privacy preserving XAI: End-users are an integral and inseparable component of XAI as they directly engage and draw insights from the content generated by these systems. While current XAI methods are predominantly model-centric, focussing on model development, evaluation and audit processes (kaplanUnifiedPracticalUsercentric2024), there is an increasing need for a shift towards a user-centric approach. When considering privacy as a system requirement or using existing privacy preserving solutions with explainability methods, human users are absent from the design and development of these approaches.
@@
  • Lack of standardised approach for privacy evaluation of XAI: Existing XAI literature provides evaluation metrics to assess aspects such as faithfulness, complexity, localisation, randomisation and robustness (hedstromQuantusExplainableAI2023). However, though privacy is fundamental in Trustworthy AI, currently there is a lack of standardised approach to evaluate the privacy of explanations. The lack of suitable quantitative metrics in this area prevents an evaluaton of safety with respect to privacy or comparison between methods for selection of those that satisfy the safety requirements of an application domain.
@@
  • Trade-offs in privacy, explainability and utility: The introduction of privacy enhancing technologies often result in adverse effects on model utility (ezzeddineDifferentialPrivacyAnomaly2024; harderInterpretableDifferentiallyPrivate2020) and the quality of explanations. For instance, the perturbation of classifier weights of support vector machines for privacy preservation is observed to deteriorate the classification accuracy and credibility of counterfactuals (mochaourabDemonstratorCounterfactualExplanations2023). Similarly, the use of differential privacy in neural network models is found to lower its accuracy (blanco-justiciaCriticalReviewUse2023) and explanation quality (liuPleaseTellMe2024). Thus when targeting these 3 properties simultaneously in an XAI system, there are often trade-offs involved.
@@
  • Lack of XAI methods that are privacy preserving by design: As emphasized by hoepmanPrivacyDesignStrategies2014, privacy is a core property of computer systems that requires addressing from system design phase, rather than treated as an add-on. However, current explainability approaches do not consider privacy as a system requirement. Due to the overlook of this safety aspect, design flaws may be introduced which can lead to unintentional privacy leakage. Adversaries may exploit these flaws to cause intentional leakage.
@@
  • Lack of privacy preserving XAI for Gen-AI and LLMs: XAI research has mainly focused on discriminative models that produce decision boundaries, and there has been limited work on developing explainability methods for Gen-AI and LLMs (schneiderExplainableGenerativeAI2024; sunInvestigatingExplainabilityGenerative2022; weiszGeneralDesignPrinciples2023). Due to the complex structure and vast number of parameters in these models, traditional explainability methods become impractical to them (zhaoExplainabilityLargeLanguage2024). These models have privacy issues, such as memorization of training data that escalates as the models become larger (carliniExtractingTrainingData2021). In addition, downstream private datasets used for in-context learning in LLMs, are found to be susceptible to membership inference (duanPrivacyRiskIncontext2023). When creating explainability approaches for these systems, privacy should be considered as an integral requirement.
@@ 7.4 Recommendations for future work
  • Improving usability of privacy preserving XAI: Explanations should be designed to meet the diverse information needs of users and integrate user-centric design principles (aliExplainableArtificialIntelligence2023; kaplanUnifiedPracticalUsercentric2024). The delivery of explanations is required to be in a format that is accessible and meaningful, taking into consideration the varying levels of expertise and requirements of different user groups. This transition would prioritize providing need-to-know information to end-users based on their specific roles within the system. Furthermore, to enhance the effectiveness of explanations, appropriate tools such as interactivity and visualization should be used to enhance the process of explanation and deepen users’ understanding (boIncrementalXAIMemorable2024). In addition, application of the 3C-principle of context, content, and consent (brunotteContextContentConsent2023) can improve the usability of XAI tools by satisfying their requirements and expectations.
@@
  • Balancing trade-off in privacy, explainability and utility: Determining the appropriate trade-off between the triad of privacy, explainability and utility, can help to achieve the right measure of balance between these properties. This could be achieved with the help of tools or tuning mechanisms. For instance, the use of compatibility matrix (abbasiFurtherInsightsBalancing2024) or hyperparameters, such as the privacy budget $\epsilon$ in differential privacy or the parameter k in k-anonymity, is useful in tuning the desired level of these required properties. Similar tuning mechanisms can be integrated in other privacy preserving approaches to achieve the required trade-off. Metrics, such as trade-off score (abbasiFurtherInsightsBalancing2024), could be useful to quantify and monitor the balance of these properties enabling researchers and practitioners to adjust the parameters based on the trade-offs involved.
@@
  • Development of XAI methods that are privacy preserving by design: The characteristics of privacy preserving XAI outlined in this review, can aid researchers and developers in building explainability methods that are privacy-enhanced by design (bozorgpanahExplainableMachineLearning2024). Furthermore, there is growing interest in neuro-symbolic approaches (hitzlerNeurosymbolicApproachesArtificial2022) and semantic technologies (seeligerSemanticWebTechnologies2019) as potential solutions as explainable-by-design strategies. Researchers can investigate the privacy safety of these techniques and integrate approaches to prioritize privacy.
  • Privacy preserving XAI for Gen-AI and LLMs: XAI plays a vital role in fostering trustworthiness (wangRationalityExplanationHuman2024) and ensuring ethical applications of Gen-AI and LLMs (luoUnderstandingUtilizationSurvey2024). However, as explainability is introduced in these models, it is important to ensure that it does not exacerbate the inherent privacy issues of these models or create new vulnerabilities. A privacy analysis of explainability methods in the early stages of development and the use of privacy attacks for auditing (carliniExtractingTrainingData2021), can boost the development of privacy-enhanced systems. Recently, Chain-of-Thought (CoT) (wei2022chain) prompting has gained traction in eliciting the step-by-step reasoning of models and providing self-explanations in the process. However, the interaction with untrusted LLMs can be a threat to users’ privacy, hence future work can be directed towards privacy preserved mechanisms (bae2025privacy) in these models. Methods such as retrieval-augmented generation (RAG) for fine tuning of outputs by augmenting external data sources (zengGoodBadExploring2024) and mechanistic interpretability (Section 2.3.8) can also be investigated as possible solutions towards achieving privacy preservation XAI. Thus, with the growing accessibility and widespread use of Gen-AI and LLMs, developing appropriate user-centric privacy preserving explainability techniques is an important avenue for further research.